OpenStudy (anonymous):

Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers. ^3sqrt(72 * t^8)

OpenStudy (anonymous):

\[\sqrt[3]{72t^8}=\sqrt[3]{2^33^2t^3t^3t^2}=2t^2\sqrt[3]{3^2t^2}\]